OpenStudy (anonymous):

Dan has $60.00 in his wallet and wants to buy new CDs. If the CDs are $11.00 each what number of CDs, x, can he buy? I need to write and solve the equation. Will medal!

OpenStudy (michele_laino):

please, if I callwith n number of purchased CDs, then I can write: \[60=n*11+r\] where r is the remainder, so you have to find the highest value of n such that r is positive

OpenStudy (anonymous):

55 is the highest he can go

OpenStudy (michele_laino):

ok! so n=

OpenStudy (anonymous):

n = 55

OpenStudy (michele_laino):

no, sorry, n=5 because 5*11=55

OpenStudy (anonymous):

Oh yes! sorry

OpenStudy (michele_laino):

ad r=?

OpenStudy (michele_laino):

oops.. and r=?

OpenStudy (anonymous):

6?

OpenStudy (michele_laino):

please, note that 55+r= 60, so r=?

OpenStudy (anonymous):

I'm working with inequalities

OpenStudy (anonymous):

5

OpenStudy (anonymous):

r=5

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

I need to write an inequality equation, will this work?

OpenStudy (michele_laino):

I think your inequliyis: \[60>11*n\] where n=5

OpenStudy (anonymous):

Oh! ok

OpenStudy (michele_laino):

oops your inequality is....

OpenStudy (anonymous):

Thank You!

OpenStudy (michele_laino):

Thank you!

OpenStudy (anonymous):

You're welcome

OpenStudy (michele_laino):

Thanks!:)